find all solutions of the equation sec4(theta)-2=0
find the solutions in the interval [0,2pi]

Answer by lwsshak3(11628)

find all solutions of the equation sec4(theta)-2=0
find the solutions in the interval [0,2pi]
**
sec4x-2=0
1/cos4x=2
cos4x=1/2
4x=π/3 and 5π/3 (in quadrants I and IV where cos>0)
x=π/12 and 5π/12